# Statistical Principles for Platform Trials

### v4

@@ 4 A Confident Platform Random Sizer App
An app [^2] Can be found at https://eouya001.shinyapps.io/ConfidentPlatformRandomSizer/ has been developed to not only illustrate the impact of conditioning on $\hat{\mu}_{SC}$ for the distribution of $V^{*}$, but also to provide insights into the calculated standard errors and unconditional correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$. We simulate a case study where we test the efficacy of five drugs. The total resources available allow for 600 people. With this setup, we can adjust the input values using the app’s sliders. Notably, ALS has a population standard deviation of 6.5 ($\sigma$=6.5), which the app uses internally to calculate standard errors of $\hat{\mu}_{SC},\hat{\mu}_{i}$ and $\hat{\mu}_{i}-\hat{\mu}_{SC}$, $i=1,\ldots,n\in N$, where $n$ is the maximum number of treatment arms erroneously approved. An example case in which sponsors should be concerned is presented in App: User Slider Controls.
@@
App: User Slider Controls $\alpha$ for Each Rx Arm: ERFw $\alpha$ for each treatment arm. Note that ERpF is $n\alpha$
@@
Standardized $\hat{\mu}_{SC}$: The shared control sample mean $\hat{\mu}_{SC}$, standardized by the standard error of the shared control sample mean
An example of an overachieving $\hat{\mu}_{SC}$, i.e. $\hat{\mu}_{SC}>0$. Different shared control values would affect the number of erroneous approvals
Maximum Number of Rx Arms Erroneously Approved: $n$, the number of treatment arms, when they are non-efficacious.
n=5
Total Sample Size: Sum of sample sizes from all treatment arms and the shared control
@@
Unconditional Correlation Between New Treatment and Control Estimates: Correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$
$\rho_{ij}=0.5$
[Uncaptioned image] Sample Size Ratio: Treatment to control sample size ratio for sample allocation. Note that while the app allows for one to go higher, it is generally not advised to go higher than $k:1$.
5:1 treatment-to-control sample size ratio
By selecting the sample size ratio, the app automatically calculates the unconditional correlation between the new treatment arm and control estimates. This feature is particularly beneficial to regulators, who have a vested interest in managing this aspect to stabilize the distribution of $V^{*}$. A lower correlation achieved by reducing the sample size ratio also decreases the variation of $V^{*}$.
@@
With an overachieving $\hat{\mu}_{SC}$, we see that the probability of erroneously approving a treatment arm is low
@@
[Uncaptioned image] Standard Errors: Gives standard errors for estimated control mean, treatment arm means, and the mean difference between control and treatment arm
@@
The app can be instrumental in designing new studies. It facilitates considerations regarding the stability of the estimates from shared controls and the (unconditional) correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$. Specifically, if regulators set a required correlation threshold, sponsors can use the app to determine the maximum sample ratio randomization that meets this criterion. The app thus serves as a valuable tool for identifying the optimal balance that satisfies all stakeholders involved.
Within a platform trial, the randomization ratio $n_{Rx}:n_{C}$ of the new compound to the shared control can vary. Ratios range from a low of $1:1$ resulting in a sample size ratio of the shared control to each new compound of $k:1$—to a high of $k:1$, where the sample size ratio reverses to $1:1$. Given a fixed total sample size for the platform trial, selecting the appropriate randomization ratio involves considering multiple factors.
Patients generally prefer a higher $n_{Rx}:n_{C}$ ratio, as it increases their chances of receiving a new treatment compound. For individual drug sponsors, the advantage of participating in a platform trial includes being able to pool control patients, which allows for a higher randomization ratio than the standard $1:1$ typically seen in standalone studies. This higher ratio can facilitate more accurate effect estimations for new compounds.
However, the perspective of the master protocol sponsor tends to favor a lower ratio, which increases the sample size of the shared control, reducing the likelihood of an overachieving control sample that complicates the demonstration of efficacy across new compounds.
From a regulatory standpoint, the stability of the ERpF control is influenced by both the estimates of the shared control’s sample means and the correlation among the pivotal statistics $T^{*}$. A lower randomization ratio enhances the sample size of the shared control, mitigating the risk of an under-performing control sample that could erroneously ease the approval of ineffective compounds. Additionally, a lower ratio reduces the correlation among pivotal statistics, thereby enhancing the stability of ERpF control.
For example, in the HEALEY ALS Platform Trial, none of the four drugs tested in Regimens ABCD demonstrated efficacy in the primary endpoint, ALSFRS-R. [^3] Active treatments for Regimen A were Zilucoplan, B was Verdiperstat, C was CNM-Au8, and D was Pridopidine. The lack of efficacy could be attributed to either the inherent ineffectiveness of the compounds, an overachieving shared control, or a combination of both. Conversely, as illustrated in App: Regulator Concerns, with an under-performed shared control sample mean had (i.e., $\hat{\mu}_{SC}<0$), there might have been a higher chance of the drugs being approved too readily, as exemplified by a $\hat{\mu}_{SC}^{\star}=-1.5$. Such a scenario would raise significant concerns for regulators.
In sum, choice of the randomization ratio $n_{Rx}:n_{C}$ should take all these considerations into account. Our app lets one find the sweet spot.
App: Regulators Concerns We now present a scenario using our app where sponsors, such as MASS General, need not be concerned, but which should capture the attention of regulators. The following inputs and outputs illustrate a hypothetical situation: $\alpha$ for Each Rx Arm: $0.05$ Standardized $\hat{\mu}_{SC}$: $-1.5$ Maximum Number of Rx Arms Erroneously Approved: $n=5$ Total Sample Size: 600 Unconditional Correlation Between New Treatments and Control estimates: 0.5 Sample Size Ratio: 5:1
[Uncaptioned image] The probability distribution of $V^{*}$ indicates that an under-performing standardized $\hat{\mu}_{SC}$ significantly increases the likelihood of erroneously approving at least one treatment, compared to its true efficacy. Regulators should be particularly concerned about this, as it suggests a non-negligible probability that ineffective drugs could mistakenly reach the market.
@@ 5.1 FDR would change policy
False Discovery Rate (FDR) control has been proposed as a method for controlling the Type I error rate across studies in Platform trials (Robertson et al., (2023)). FDR is defined as the expected proportion of incorrect rejections (rejecting a true null hypothesis) among all rejections, both correct and incorrect  [^4] Rejecting a false null hypothesis is considered “correct”, while rejecting a true null hypothesis is “incorrect”..
@@
$\mathrm{FDR}$ $=$ $\mathrm{E}\left[\frac{\mbox{number of rejected }H_{0i}\mbox{ thatare true}}{\mbox{total number of rejected }H_{0i}}\right]$
$=$ $\mathrm{E}\left[\frac{\mbox{number of in-efficacious compounds incorrectly approved }}{\mbox{total number of compounds approved, both efficacious and in-efficacious}}\right]$
While current regulatory practice of controlling FAR (which is Tukey’s per family error rate across studies) limits the number of in-efficacious compounds incorrectly approved to 5-per-100, controlling FDR in Platform trials would increase that number to beyond 5-per-100 as we will show below.
Recall $N$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy and $|N|=n$. For each $i\in N$, $Z_{i}$ is the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)$, and
$V^{*}=\sum_{i\in N}Z_{i}.$
@@
$\dfrac{\mathrm{EFpF}}{n}$ $=$ $\dfrac{E(V^{*})}{n}$ (5.2)
$=$ $\dfrac{\sum_{i=1}^{n}P\left\{Z_{i}=1\right\}}{n}$ (5.3)
$=$ $\sum_{v=1}^{n}\dfrac{v}{n}P\{V^{*}=v\}$ (5.4)
@@
$\mathrm{FDR}$ $=$ $\sum_{v=1}^{n}\dfrac{v}{v+n^{\mathrm{false}}}P\{V^{*}=v\}$ (5.5)
$\leq$ $\sum_{v=1}^{n}\dfrac{v}{0+n^{\mathrm{false}}}P\{V^{*}=v\}.$ (5.6)
@@
So, if there is a sequence of many Platform trials with $k=2$ studies following this scenario, then by the LLN (2.1), the False Approval Rate (FAR) is 10-per-100 (not 5-per-100).
Switching to FDR control would thus be a policy change: For every efficacious compound approved, regulators reward the industry with an additional fungible False Approval Rate (FAR) token [^5] We term the tokens fungible because any company can cash them in for any compound. toward approval of in-efficacious compounds.
Adopting FDR control would represent a significant policy shift. Under this system, for every efficacious compound approved, regulators would essentially reward the industry with an additional fungible False Approval Rate (FAR) token [^6] We describe the tokens as fungible because any company can use them to seek approval for any compound. that could be used toward the approval of inefficacious compounds. It is important to note that we find no explicit language in the Food, Drug, &\&& Cosmetic Act or the Kefauver-Harris Amendment suggesting that the number of efficacious compounds correctly approved affects the guarding of the number of in-efficacious compounds incorrectly approved.
@@ 5.2 Integrity of error rates against manipulation is desirable
Whereas the False Approval Rate (FAR) is only concerned with the rate at which in-efficacious compounds are incorrectly approved, clearly the issue with FDR is its denominator contains both correct and incorrect approvals of compounds. Similar to the concern in Section 6 of Finner and Roter (2001), we wonder whether that might subject FDR to manipulation as follows.

### v5

@@ 3.3.3 A Confident Platform Random Sizer (Interactive visualization via Shiny App)
In a simulated case study testing the efficacy of five drugs, the total resources allows for the participation of 600 people. With this setup, users can adjust the input values using the app’s sliders. Notably, ALS has a population standard deviation of 6.5 ($\sigma$ = 6.5), which the app uses internally to calculate standard errors of $\hat{\mu}_{SC},\hat{\mu}_{i}$ and $\hat{\mu}_{i}-\hat{\mu}_{SC}$, $i=1,\ldots,k$, where $k$ is the number of treatment arms. A scenario of particular concern to sponsors is detailed in App: User Slider Controls (Figure 2)
By selecting the sample size ratio, the app automatically calculates the unconditional correlation between the new treatment arm and control estimates. This feature is particularly beneficial to regulators, who have a vested interest in managing this aspect to stabilize the distribution of $V^{*}$. A lower correlation achieved by reducing the sample size ratio also decreases the variation of $V^{*}$. Therefore, the app can be instrumental in designing new studies. If regulators set a required correlation threshold, sponsors can use the app to determine the optimal sample size ratio randomization that meets this criterion. Based on the sample size allocation, the distribution of $V^{*}$ conditional on the sample mean of the shared control can also be evaluated (Figure 3).
Within a platform trial, the randomization ratio $n_{Rx}:n_{C}$ of the new compound to the shared control can vary from a low of $1:1$ - resulting in a sample size ratio of the shared control to each new compound of k:1k:1—to a high of $k:1$, where the ratio reverses to $1:1$. Given a fixed total sample size for the platform trial, selecting the appropriate randomization ratio involves multiple factors. As shown in Chandereng et al (2020), if minimizing the variance of each individual treatment vs. control $Rx:C$ comparison were the sole consideration, then as described in Appendix A of FDA (2023), one would set $n_{Rx}:n_{C}$ to be essentially $\sqrt{k}$. This ratio can be a starting point for all the stakeholders to explore the choice of $n_{Rx}:n_{C}$.
Patients generally prefer a higher $n_{Rx}:n_{C}$ ratio, as it increases their chances of receiving a new treatment compound. For individual drug sponsors, the advantage of participating in a platform trial includes being able to pool control patients, which allows for a higher randomization ratio than the standard $1:1$ typically seen in stand-alone studies. This higher ratio can facilitate more accurate effect estimations for new compounds.
However, the perspective of the master protocol sponsor tends to favor a lower ratio, which increases the sample size of the shared control, reducing the likelihood of an over-achieving control sample that complicates the demonstration of efficacy across new compounds.
From a regulatory standpoint, the stability of the IAR control is influenced by both the shared control’s sample means and the correlation among the pivotal statistics $T^{*}$. A lower ratio reduces the correlation among pivotal statistics, thereby enhancing the stability of IRA control.
In the HEALEY ALS Platform Trial, none of the four drugs tested in Regimens ABCD demonstrated efficacy in the primary endpoint, ALSFRS-RdddActive treatments for Regimen A were Zilucoplan, B was Verdiperstat, C was CNM-Au8, and D was Pridopidine.. The lack of efficacy could be attributed to either the inherent ineffectiveness of the compounds, an over-achieving shared control, or a combination of both. Conversely, as illustrated in App: Regulators Concerns, with an under-performing shared control sample mean there might have been a higher chance of the drugs being approved too readily, as exemplified by a $\hat{\mu}_{SC}$ that is 1.5 standard deviations below the $\mu_{SC}$. Such a scenario would raise significant concerns for regulators (Figure 4).
In sum, choice of the randomization ratio $n_{Rx}:n_{C}$ should take all these into account. Our app lets one find the sweet spot.
@@
App: User Slider Controls $\alpha$ of Each Rx Arm: ERFw $\boldsymbol{\alpha}$ for each treatment arm. Note that ERpF is $k\alpha$
@@
$\boldsymbol{\delta:\mu_{SC}\pm\delta\sigma_{SC}}$: how many standard deviations is $\hat{\mu}_{SC}$ away from its true mean. An example of 0.5 SD away from $\mu_{SC}$ (an over-achieving $\hat{\mu}_{SC}$) Number of Non-Efficacious Rx Arms: $\ell$, assuming all treatment arms $k$ are non-efficacious, $l=k$. An example of $l=k=5$ Total Sample Size: sample sizes including all treatment arms and the shared control
@@
Correlation: $\rho_{ij}$, Correlation between $T_{i}^{\star}$ and $T_{j}^{\star}$
An example of $\rho_{ij}=0.5$
Rx Arm: Control Sample Size Ratio: Treatment to control sample size ratio. Note that the app allows for one to go higher, but we advise not to go higher than $k:1$.
An example of 5:1 treatment-to-control sample size ratio
[Uncaptioned image] Figure 2: User Slider Controls
@@
With an over-achieving $\hat{\mu}_{SC}$, we see that the probability of erroneously approving a treatment arm is low
@@
[Uncaptioned image] Figure 3: App Output Standard Errors: Gives standard errors for estimated control mean, treatment arm means, and the difference between the sample mean of the shared control and that of the treatment arm
@@
App: Regulators Concerns We now present a scenario using our app where sponsors, such as MASS General, need not be concerned, but which should capture the attention of regulators. The following inputs and outputs illustrate a hypothetical situation: Slider Inputs $\boldsymbol{\alpha}$ for Each Rx Arm: $0.05$ $\boldsymbol{\delta}$: 1.5-1.5 Number of Non-Efficacious Rx Arms: $\ell=5$ Total Sample Size: 600 Correlation $\boldsymbol{\rho_{ij}}$: 0.5 Sample Size Ratio: 5:1
[Uncaptioned image] Figure 4: App Output The probability distribution of $V^{*}$ indicates that an under-performing $\hat{\mu}_{SC}$ significantly increases the likelihood of erroneously approving at least one treatment, compared to its true efficacy. Regulators should be particularly concerned about this, as it suggests a non-negligible probability that ineffective drugs could mistakenly reach the market.
@@ 4.1 FDR would change policy
False Discovery Rate (FDR) control has been proposed as a method for controlling the Type I error rate across studies in Platform trials (Robertson et al.,, 2023). FDR is defined as the expected proportion of incorrect rejections (rejecting a true null hypothesis) among all rejections, both correct and incorrecteeeRejecting a false null hypothesis is considered “correct”, while rejecting a true null hypothesis is “incorrect”..
@@
[…]
While current regulatory practice of controlling IAR (ERpF) across studies limits the number of in-efficacious compounds incorrectly approved to 5-per-100, controlling FDR in Platform trials would increase that number to beyond 5-per-100 as we will show below.
Recall $L$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy and $|L|=l$. For each $i\in L$, $Z_{i}$ is the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)$, and
$V^{k}=\sum_{i\in L}Z_{i}.$
@@
$\dfrac{\mathrm{EFpF}}{l}$ $=$ $\dfrac{E(V^{k})}{l}$ (4.2)
$=$ $\dfrac{\sum_{i=1}^{l}P\left\{Z_{i}=1\right\}}{l}$ (4.3)
$=$ $\sum_{v=1}^{l}\dfrac{v}{l}P\{V^{k}=v\}$ (4.4)
@@
$\mathrm{FDR}$ $=$ $\sum_{v=1}^{l}\dfrac{v}{v+n^{\mathrm{false}}}P\{V^{k}=v\}$ (4.5)
$\leq$ $\sum_{v=1}^{l}\dfrac{v}{0+n^{\mathrm{false}}}P\{V^{k}=v\}.$ (4.6)
@@
So, if there is a sequence of many Platform trials with $k=2$ studies following this scenario, then by the LLN (2.1), the Incorrect Approval Rate (IAR) is 10-per-100 (not 5-per-100).
Adopting FDR control would represent a significant policy shift. Under this system, for every efficacious compound approved, regulators would essentially reward the industry with an additional fungible False Approval Rate (FAR) tokenfffWe describe the tokens as fungible because any company can use them to seek approval for any compound. that could be used toward the approval of inefficacious compounds. It is important to note that we find no explicit language in the Food, Drug, &\& Cosmetic Act or the Kefauver-Harris Amendment suggesting that the number of efficacious compounds correctly approved affects the guarding of the number of in-efficacious compounds incorrectly approved.
@@ 4.2 Integrity of error rates against manipulation is desirable
Whereas the IAR is only concerned with the rate at which in-efficacious compounds are incorrectly approved, clearly the issue with FDR is its denominator contains both correct and incorrect approvals of compounds. Similar to the concern in Section 6 of Finner and Roter(2001), we wonder whether that might subject FDR to manipulation as follows.
@@
4.3 Struggles of FDR Not Restricted to Platform Trials
FDR control also encounters serious limitations in other settings. In large-scale Genome-Wide Association Studies (GWAS), each hypothesis typically tests whether a SNP has no association with the disease outcome. Classical FDR methods assume that most hypotheses are truly null because then controlling FDR limits the number of erroneous findings among the reported results. However, as demonstrated in Ding et al., (2018) and further elaborated in Ding et al., (2021), due to linkage disequilibrium (LD), SNPs in close proximity are highly correlated. Consequently, when a single causal SNP influences a biological pathway, many neighboring SNPs may also appear statistically significantly associated with the disease outcome. In such settings, the complete null hypothesis may be false for a large proportion of SNPs, leading to a so-called “zero-null” scenario, and further resulting in small number of false discoveries $V$ relative to the total number of rejections $R$, and therefore
$\mathrm{FDR}=\mathbb{E}\left[\frac{V}{R}\right]\approx 0$
automatically. Obviously, controlling FDR doesn’t tell us much about the quality of discoveries and a coherent alternative was proposed as a per-panel error rate (Ding et al.,, 2018, 2021).
4.3.1 Crosstalk plots: A Visual Tool for GWAS SNP Behavior
To further illustrate why FDR becomes difficult to interpret in GWAS, we introduce crosstalk plots as a geometric representation of SNP dependence. To construct a crosstalk plot, select two SNPs, designating one as the conditioning SNP, $C$ and the other as the response or target SNP, $T$. This designation is purely analytical and does not imply causality. The sample is partitioned according to the genotype of $C$, where $C\in\{0,1,2\}$ denote the number of minor alleles i.e., homozygous major, heterozygous, and homozygous minor. These three groups form the layers of a crosstalk plot.
Within each stratum of $C$, we compute the conditional genotype distribution of the three genotypes of $T$. Each layer has a probability vector that sums to one, allowing it to be represented as a barycentric point on one of the three planes corresponding to the genotype strata of $C$. Connecting the three barycentric points across layers yields a trajectory whose geometry reflects dependence between the two SNPs. Figure 5(a) and 5(b) provide toy examples illustrating how the plot can demonstrate independence and dependence appearance and can be translated into a 2D projection form, similar to that seen in Ding et al., (2021).
(a) Side view
(b) Top-down view
Figure 5: Plotting of three different toy examples. Panel (a) depicts the side view, where vertical alignment corresponds to geometric orthogonality across ternary layers, and panel (b) depicts the corresponding top-down projection used in the original two-dimensional plot. Each colored trajectory connects the conditional genotype distribution of a target SNP across the three genotype strata of a conditioning SNP, which is able to be plotted on stacked ternary planes corresponding to $C=0,1,2$ due to them summing up to 1. The black trajectory represents an independent SNP; by having identical conditional probabilities for each strata connecting each point forms a line orthogonal to each ternary plane that collapses to a single point under top-down projection in (b). As such, a single point in the projected view corresponds geometrically to orthogonality and probabilistically to independence. The blue trajectory depicts mild dependence, where connecting the small shifts across strata with a line produces a smaller V-shape in the projected view. The red trajectory shows strong dependence, where the connection between the large shifts in conditional probabilities across the strata generates a pronounced V-shape.
A single dependent SNP produces one V-shaped trajectory under projection; when multiple SNPs are displayed relative to the same causal SNP, these individual V-shaped trajectories accumulate into larger geometric patterns. This is one of the advantages of crosstalk plots over many existing SNP visualization methods as they allow multiple SNPs to be layered within the same geometric framework. For example, a mosaic plot is limited to displaying a single contingency table at a time, so each SNP pair must be examined separately. In contrast, a crosstalk plot represents each SNP as a trajectory of conditional genotype probabilities across strata of a conditioning SNP, allowing many SNPs to be displayed simultaneously within one ternary coordinate system. This makes it possible to compare the behavior of all SNPs relative to the same conditioning structure and to detect common patterns that would otherwise be difficult to recognize across many separate plots. In Figure 6, we demonstrate this using the apolipoprotein E (APOE) locus, which has been central to Alzheimer disease research for decades. Variation at this locus is defined by two key SNPs, rs429358 and rs7412, which give rise to three common APOE isoforms (APOE2, APOE3, APOE4) (Belloy et al.,, 2019). Since these variants jointly determine APOE isoforms, they provide a natural case study for Crosstalk plots.
Figure 6: When many SNPs are layered together, shared dependence patterns become visually apparent that are not evident when examining individual SNP pairs. Using reference data from The 1000 Genomes Project Consortium, (2015), we designate the apolipoprotein E SNP rs429358 as the causal SNP on which multiple SNPs are conditioned on. If neighboring markers were independent of the causal SNP, most projected trajectories would collapse to isolated points. However, the trajectories form repeated V-shapes, following a common directional trend and leaving a characteristic curved white space due to constraints imposed by the Hardy-Weinberg equilibrium (that the proportion of AaAa is $2pq$ with $P\{A\}=p$ and $P\{a\}=q$) together with dependence induced by linkage disequilibrium, indicating that the SNPs are not behaving independently but are shifting systematically across genotype strata. This geometric clustering has direct inferential implications in GWAS. Once one SNP is truly associated with an outcome, many dependent SNPs in the same region also become statistically associated because they carry partial statistical information about the same underlying signal as previously noted by Ding et al., (2021). As a result, large numbers of markers may appear significant even though they do not represent distinct biological discoveries. In such settings, false discovery rate (FDR) becomes difficult to interpret: the number of rejected hypotheses may become very large, not because many independent signals are present, but because one dependence structure generates repeated manifestations of the same association. Crosstalk plots make this phenomenon visible by showing that many significant SNPs often trace the same geometric pattern, reflecting underlying dependence structure rather than representing separate independent effects, underscoring why FDR can lose interpretability in GWAS.
5 Discussion
While the applications of the conditionality principle and layering of error rates were focused on platform trials, there are numerous applications elsewhere. In fact, the effects of implicit conditioning are already evident in regulatory practice and drug labeling. A drug label is only constructed after an approval decision has been made. The act of labeling therefore occurs after conditioning on primary endpoint success, and the drug label represents a summary of evidence given that condition; additional claims must be supported by inference that is valid given primary end point success. Accordingly, regulators do not ask whether secondary endpoints perform marginally across all trials, but only in the paths that have achieved success on the primary endpoint.
Similarly, the layering of error rates naturally arises because decisions are made both at the trial level and across the broader regulatory landscape. Within a specific study, incorrect conclusions about any endpoint may lead to erroneous claims, requiring strict control. We therefore want to ensure that the ERFw is controlled, i.e. the probability of making at least one incorrect inference among the endpoints in the trial does not exceed the prespecified level. At the higher layer, regulators also desire control of the error rate with respect to the approval of a drug compound, corresponding to IAR.
We also note that crosstalk plots have additional uses outside those listed in this paper. Such usages include visualizing linkage disequilibrium of a SNP.